OpenStudy (anonymous):

Which of the following expressions shows how to rewrite 7 − 8 using the additive inverse and displays the expression correctly on a number line? @mathstudent55

OpenStudy (mathstudent55):

Do you know what the additive inverse is?

OpenStudy (anonymous):

um not really :( srry

OpenStudy (mathstudent55):

Additive inverses are numbers that add up to zero.

OpenStudy (anonymous):

oh ok

OpenStudy (mathstudent55):

That is the definition. Here is the practical way to know it. The additive inverse of zero is zero since 0 + 0 = 0

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

With all other numbers, the additive inverse of a number is just take the number and change the sign.

OpenStudy (mathstudent55):

The additive inverse of 2 is -2. The additive inverse of -1/2 is 1/2 The additive inverse of 7.18 is -7.18

OpenStudy (mathstudent55):

Ok so far?

OpenStudy (anonymous):

yes:)

OpenStudy (mathstudent55):

Good. When you deal with subtraction, you can define subtraction as adding the additive inverse.

OpenStudy (mathstudent55):

That means any subtraction can be treated as an addition, but what you must do is change the number you are subtracting into its additive inverse.

OpenStudy (mathstudent55):

For example, change 8 - 2 into an addition. Answer: 8 - 2 = 8 + (-2)

OpenStudy (anonymous):

oh ok

OpenStudy (mathstudent55):

The example above was the subtraction 8 - 2. The number 2 needs to be subtracted from the number 8. We change the subtraction into an addition by changing the minus sign into a plus sign AND we change the 2 into the additive inverse of 2, which is -2.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

Now look at your example. You have the subtraction 7 - 8 To change the subtraction into an addition, you keep the 7. Change the - to + Then change the 7 to its additive inverse which is -7

OpenStudy (mathstudent55):

Try it and show me what you get.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ummm... is it -8+7?

OpenStudy (anonymous):

is that the answer?

OpenStudy (anonymous):

helloooooo???

OpenStudy (mathstudent55):

7 - 8 Keep the 7 as is: 7 Change the - into a + 7 + Chang ethe 8 into its additive inverse: -8 7 + (-8)

OpenStudy (mathstudent55):

7 + (-8) is the same as -8 + 7

OpenStudy (mathstudent55):

Do you have choices?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh wait no

OpenStudy (anonymous):

so whats the final answer?

OpenStudy (mathstudent55):

I say it's 7 + (-8)

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

can you help me with more?

OpenStudy (mathstudent55):

yes, one more, then i have to go.

OpenStudy (anonymous):

ok